Question 1 (40 marks)

Magnate Ltd is a manufacturing company which produces a fixed budget for planning purposes. Set out below is the monthly fixed budget of production costs, together with the actual results observed for the month of November 2019.

Budget

Actual

Units produced

5,000

5,500

£

£

Cost

Direct material

20,000

22,764

Direct labour

60,000

75,900

Variable Production overhead

14,000

14,950

Fixed Production overhead

10,000

9,000

Depreciation

4,000

4,000


In preparing the flexed budget, the following standards were adopted:

Direct material                       10 kg of materials per unit produced Direct labour    2 hours per unit produced


Variable overheads               Allocated on the basis of direct labour hours

The following additional information is available concerning the actual output:

  1. The actual usage of materials in November was 54,200 kg: and
  2. The wage rate increased to £6.60 per hour at the start of November.

Required:

  1. Show in tabular form the original budget, and the flexed and actual budget with variances in respect of the production costs for the month of November 2019.

(12 marks)

  1. Calculate all relevant cost variances in as much detail as possible.

(24 marks)

  1. Suggest possible reasons for variances in the question (b). Why management may wish to investigate these variances.

(4 marks) (Question 1 total: 40 marks)

Question 1 (40 marks)

Magnate Ltd is a manufacturing company which produces a fixed budget for planning purposes. Set out below is the monthly fixed budget of production costs, together with the actual results observed for the month of November 2019.

Budget

Actual

Units produced

5,000

5,500

£

£

Cost

Direct material

20,000

22,764

Direct labour

60,000

75,900

Variable Production overhead

14,000

14,950

Fixed Production overhead

10,000

9,000

Depreciation

4,000

4,000


In preparing the flexed budget, the following standards were adopted:

Direct material                       10 kg of materials per unit produced Direct labour    2 hours per unit produced


Variable overheads               Allocated on the basis of direct labour hours

The following additional information is available concerning the actual output:

  1. The actual usage of materials in November was 54,200 kg: and
  2. The wage rate increased to £6.60 per hour at the start of November.

Required:

  1. Show in tabular form the original budget, and the flexed and actual budget with variances in respect of the production costs for the month of November 2019.

(12 marks)

  1. Calculate all relevant cost variances in as much detail as possible.

(24 marks)

  1. Suggest possible reasons for variances in the question (b). Why management may wish to investigate these variances.

(4 marks) (Question 1 total: 40 marks)
Answer #1

Ans.(a).

Original, Flexed and Actual
Original Flexed Actual
Units Produced 5,000 5,500 5,500
£ £ £
Cost
Direct Material 20,000 22,000 22,764
Direct Labour 60,000 66,000 75,900
Variable Production Overhead 14,000 15,400 14,950
Fixed Production Overheads 10,000 10,000 9,000
Depreciation 4,000 4,000 4,000

Ans.(b). Material Cost Variance = 764 (A)

Material Price Variance = (0.4 - 0.42) x 54,200 = 1,084 (A)

Material Usage Variance = (55,000 - 54,200) x 0.40 = 320 (F)

Labour Cost Variance = 9,900 (A)

Labour Rate Variance = (6 - 6.6) x 11,500 = 6,900 (A)

Labour Efficiency Variance = (11,000 - 11,500) x 6 = 3,000 (A)

Variable Overhead Cost Variance = 450 (F)

Variable Overhead Rate Variance = (1.4 - 1.3) x 11,500 = 1,150 (F)

Variable Overhead Efficiency Variance = (11,000 - 11,500) x 1.4 = 700 (A)

Fixed Overhead Cost Variance = (2 x 5,500) - 9,000 = 2,000 (F)

Fixed Overhead Expenditure Variance = (10,000 x 1) - 9,000 = 1,000 (F)

Fixed Overhead Volume Variance = (5,500 - 5,000) x 2 = 1,000 (F)